# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ B.4 Proof of Theorem 4
$\leq$ $n_{T}\max_{j\leq n_{T}}\sum_{l=1}^{p_{j}}P\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left|\sum_{t=1}^{T}(x_{j,t}x_{j,t-l}-E(x_{j,t}x_{j,t-l}))}\right|>CT}\right).$
Now, we apply the Triplex inequality (Jiang, 2009) and obtain
$P\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left|\sum_{t=1}^{T}(x_{j,t}x_{j,t-l}-E(x_{j,t}x_{j,t-l})}\right|>CT}\right)\leq 2m\ \text{exp}\mathopen{}\mathclose{{}\left(\frac{-TC^{2}}{288m^{2}\kappa^{2}}}\right)+$
$\cfrac{6}{CT}\sum_{t=1}^{T}E\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|E(x_{j,t}x_{j,t-l}|\mathcal{F}_{t-m})-E(x_{j,t}x_{j,t-l})}\right|}\right]+$
$\cfrac{15}{CT}\sum_{t=1}^{T}E\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|x_{j,t}x_{j,t-l}}\right|1_{\{|x_{j,t}x_{j,t-l}|>\kappa\}}}\right]\coloneqq I_{1}+I_{2}+I_{3}\ \ .$
for any $\kappa>0$. For the first term we have
$n_{T}\sum_{l=1}^{p_{j}}I_{1}=2n_{T}p_{i}m\ \text{exp}\mathopen{}\mathclose{{}\left(\frac{-TC^{2}}{288m^{2}\kappa^{2}}}\right),$
so we need $n_{T}p_{i}m\ \text{exp}\mathopen{}\mathclose{{}\left(\frac{-T}{m^{2}\kappa^{2}}}\right)\rightarrow 0$.
Without loss of generality, by assumptions 1 and 3, as a consequence of the Cauchy-Schwarz inequality, and Theorems 17.5,17.8-17.10 in Davidson (1994), $\{x_{j,t}x_{j,t-l}-E(x_{j,t}x_{j,t-l})\}$ is $b_{2}$-bounded and $L_{b_{1}}$-mixingale with respect to $\mathcal{F}_{ut}\coloneqq\sigma(\mathbf{u}_{t},\mathbf{u}_{t-1},\dots)$, with non negative mixingale constants $d_{t}\leq C$ and mixingale sequence $\nu_{m}$ of size -$a$, with $a\geq 1$. Therefore, for the Jensen inequality, we have that $E\mathopen{}\mathclose{{}\left(|E(x_{j,t}x_{j,t-l}|\mathcal{F}_{t-m})-E(x_{j,t}x_{j,t-l})|}\right)\leq d_{t}\nu_{m}$. Thus, for the second term we have
$I_{2}\leq\frac{6}{CT}\sum_{t=1}^{T}d_{t}\nu_{m}\leq C\nu_{m};\ \ \ \ \ n_{T}\sum_{l=1}^{p_{j}}I_{2}\leq Cn_{T}p_{i}m^{-a},$
so we need $n_{T}p_{i}m^{-a}\rightarrow 0$.
For the third term, by Holder’s and Markov’s inequality we have that
$E\mathopen{}\mathclose{{}\left(|x_{j,t}x_{j,t-l}|1_{|\{x_{j,t}x_{j,t-l}\}|>\kappa}}\right)\leq\mathopen{}\mathclose{{}\left(E\mathopen{}\mathclose{{}\left(|x_{j,t}x_{j,t-l}|^{b_{1}}}\right)}\right)^{1/b_{1}}\mathopen{}\mathclose{{}\left(\frac{E\mathopen{}\mathclose{{}\left(|x_{j,t}x_{j,t-l}|^{b_{1}}}\right)}{\kappa^{b_{1}}}}\right)^{1-1/b_{1}}\leq$
$\leq\kappa^{1-b_{1}}E\mathopen{}\mathclose{{}\left(|x_{j,t}x_{j,t-l}|^{b_{1}}}\right),\ \ \ \ \ n_{T}\sum_{l=1}^{p_{j}}I_{3}\leq Cn_{T}p_{i}\kappa^{1-b_{1}},$
so we need $n_{T}p_{i}\kappa^{1-b_{1}}\rightarrow 0.$
We jointly bound the three terms by a sequence $\zeta_{T}\rightarrow 0$ as follows:
$Cn_{T}p_{i}m\ \text{exp}\mathopen{}\mathclose{{}\left(\frac{-T}{m^{2}\kappa^{2}}}\right)\leq\zeta_{T},\ \ \ \ \ Cn_{T}p_{i}m^{-a}\leq\zeta_{T},\ \ \ \ \ Cn_{T}p_{i}\kappa^{1-b_{1}}\leq\zeta_{T}.$
First, note that $m\in\mathbb{Z}^{+},\ \kappa>0$. Moreover, we assume that $\frac{\zeta_{T}}{n_{T}p_{i}}\leq\frac{1}{e}\Longrightarrow\frac{\zeta_{T}}{n_{T}p_{i}m}\leq\frac{1}{e}$. We isolate $\kappa$ from the first and third terms
$Cn_{T}p_{i}m\ \text{exp}\mathopen{}\mathclose{{}\left(\frac{-T}{m^{2}\kappa^{2}}}\right)\leq\zeta_{T}\ \ \Longleftrightarrow\ \ \kappa\leq C\frac{\sqrt{T}}{m}\frac{1}{\text{log}\mathopen{}\mathclose{{}\left(n_{T}p_{i}m/\zeta_{T}}\right)}\ \ ,$
$Cn_{T}p_{i}\kappa^{1-b_{1}}\leq\zeta_{T}\ \ \Longleftrightarrow\ \ \kappa\geq C(n_{T}p_{i})^{1/(b_{1}-1)}\zeta_{T}^{-1/(b_{1}-1)}\ \ .$
Since we have a lower and upper bound on $\kappa$, we need to make sure both bounds are satisfied:
$C_{1}(n_{T}p_{i})^{1/(b_{1}-1)}\zeta_{T}^{-1/(b_{1}-1)}\leq C_{1}\frac{\sqrt{T}}{m}\frac{1}{\sqrt{\text{log}\mathopen{}\mathclose{{}\left(n_{T}p_{i}m/\zeta_{T}}\right)}}$
$\Longleftrightarrow m\sqrt{\text{log}\mathopen{}\mathclose{{}\left(n_{T}p_{i}m/\zeta_{T}}\right)}\leq C\sqrt{T}(n_{T}p_{i})^{-1/(b_{1}-1)}\zeta_{T}^{1/(b_{1}-1)}\ \ .$
Isolating $m$, we have that $Cn_{T}p_{i}m^{-a}\leq\zeta_{T}\Longleftrightarrow m\geq C(n_{T}p_{i})^{1/a}\zeta_{T}^{-1/a}.$ Assuming $\frac{\zeta_{T}}{n_{T}p_{i}m}\leq\frac{1}{e}$, we have that $m\leq m\sqrt{\text{log}(n_{T}p_{i}m/\zeta_{T})}$ and therefore we need to ensure that
$C(n_{T}p_{i})^{1/a}\zeta_{T}^{-1/a}\leq C\sqrt{T}(n_{T}p_{i})^{-1/(b_{1}-1)}\zeta_{T}^{1/(b_{1}-1)}$
$\Longleftrightarrow\sqrt{T}\geq C(n_{T}p_{i})^{\frac{b_{1}+a-1}{a(b_{1}-1)}}\zeta_{T}^{-\frac{b_{1}+a-1}{a(b_{1}-1)}}\ \ .$ (B.11)
Thus, when (B.11) is satisfied, $n_{T}\sum_{l=1}^{p_{j}}I_{1}+I_{2}+I_{3}\leq 3\zeta_{T}$ and $\Pr(\mathcal{B}_{T})\geq 1-3\zeta_{T}$. $\blacksquare$
Note that under assumption 3 we have that $\Pr(\mathcal{B}_{T})\geq 1-2\zeta_{T}$, since $Cn_{T}p_{i}m^{-a}=0$, instead of $1-3\zeta_{T}$, given that $\{x_{j,t}x_{j,t-l}\}$ is a $L_{m_{1}}$-mixingale of size $\infty$- ∞.
Lemma B.6:
Let $\mathcal{A}_{T}\coloneqq\mathopen{}\mathclose{{}\left\{\underset{i\leq n_{T}+p_{y},t\leq T}{\text{max}}\mathopen{}\mathclose{{}\left|\sum_{t=1}^{T}w_{i,t}v_{t}}\right|\leq\frac{T\lambda}{4}}\right\}$. Under Assumption 1 and 4, for $\zeta_{T}$ as in Lemma B.5, if $\lambda\geq\frac{C(n_{T}+p_{y})^{1/c_{1}}\zeta_{T}^{-1/c_{1}}}{\sqrt{T}}$, then $\Pr(\mathcal{A}_{T})\geq 1-\zeta_{T}$.
Proof of Lemma B.6: By Assumptions 1, 4 and Theorems 17.5, 17.9 and 17.10 in Davidson (1994), we have that $\{w_{i,t}v_{i,t}\}$ is an $L_{m}$-mixingale of appropriate size. By the union bound, the Markov’s inequality and the Hansen’s mixingale concentration inequality, it follows that
@@
$P\mathopen{}\mathclose{{}\left(\max_{i\leq n_{T}+p_{y},l\leq T}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\sum_{t=1}^{l}w_{i,t}v_{t}}\right|}\right]>\frac{T\lambda}{4}}\right)\leq\sum_{i=1}^{n_{T}+p_{y}}P\mathopen{}\mathclose{{}\left(\max_{l\leq T}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\sum_{t=1}^{l}w_{i,t}v_{t}}\right|}\right]>\frac{T\lambda}{4}}\right)\leq$
@@
$\mathopen{}\mathclose{{}\left(\frac{T\lambda}{4}}\right)^{-c_{1}}\sum_{i=1}^{n_{T}+p_{y}}E\mathopen{}\mathclose{{}\left[\max_{l\leq T}\mathopen{}\mathclose{{}\left|\sum_{t=1}^{l}w_{i,t}v_{t}}\right|^{c_{1}}}\right]\leq\mathopen{}\mathclose{{}\left(\frac{T\lambda}{4}}\right)^{-c_{1}}\sum_{i=1}^{n_{T}+p_{y}}C_{1}^{c_{1}}\mathopen{}\mathclose{{}\left(\sum_{t=1}^{T}d_{t}^{2}}\right)^{c_{1}/2}\leq$
@@
$C(n_{T}+p_{y})T^{c_{1}/2}\mathopen{}\mathclose{{}\left(\frac{T\lambda}{4}}\right)^{-c_{1}}\ \ .$
@@
This means that $\Pr(\mathcal{A}_{T})\geq 1-C\mathopen{}\mathclose{{}\left(n_{T}+p_{y}}\right)\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{T}\lambda}}\right)^{c_{1}}$. The Lemma follows from choosing $(n_{T}+p_{y})(\lambda\sqrt{T})^{-c_{1}}\leq\zeta_{T}$, for which $\lambda\geq\frac{C(n_{T}+p_{y})^{1/c_{1}}\zeta_{T}^{-1/c_{1}}}{\sqrt{T}}$. $\blacksquare$
Proof of Theorem 4: We combine the results of Lemmas B.5 and B.6. By Lemma B.6 we have $\Pr(\mathcal{A}_{T})\geq 1-\zeta_{T}$ when $\lambda\geq C\zeta_{T}^{-1/c_{1}}\frac{(n_{T}+p_{y})^{1/c_{1}}}{\sqrt{T}}$. For Lemma B.5 we need $\zeta_{T}\leq\frac{n_{T}p_{i}}{e}$, $\sqrt{T}\geq C(n_{T}p_{i})^{\frac{b_{1}+a-1}{a(b_{1}-1)}}\zeta_{T}^{-\frac{b_{1}+a-1}{a(b_{1}-1)}}$. Therefore, for $n$ and $T$ sufficiently large, we have $\Pr(\mathcal{B}_{T})\geq 1-\zeta_{T}$. The result then follows by the union bound $\Pr(\mathcal{A}_{T}\cap\mathcal{B}_{T})\geq 1-C\zeta_{T}\rightarrow 1$ as $n_{T},T\rightarrow\infty$. The result of the Theorem follows by choosing $\zeta_{T}=\mathopen{}\mathclose{{}\left(\sqrt{\text{log(T)}}}\right)^{-1}$. $\blacksquare$
B.5 Proof Theorem 5
The proof is based on the relevant contribution on LASSO oracle inequalities provided in Chapter 6 of Bühlmann and van de Geer (2011).
By Lemma 6.1 in Bühlmann and van de Geer (2011) we obtain
@@ B.5 Proof Theorem 5
The penalty $\lambda$ is chosen such that $T^{-1}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\mathbf{W}\mathbf{v}}\right|}\right|_{\infty}\leq\lambda$. The event $\mathcal{A}_{T}\coloneqq\mathopen{}\mathclose{{}\left\{T^{-1}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\mathbf{W}\mathbf{v}}\right|}\right|_{\infty}\leq\frac{\lambda_{0}}{2}}\right\}$ has to hold with high probability, where $\lambda_{0}\leq\frac{\lambda}{2}$. Lemma B.6 and Theorem 4 of the main text prove that $\Pr(\mathcal{A}_{T})\geq 1-\zeta_{T}$ and $\Pr(\mathcal{A}_{T}\cap\mathcal{B}_{T})\geq 1-C\zeta_{T}$ for some $C>0$.
Since $\lambda\geq 2\lambda_{0}$ under $\mathcal{A}_{T}$ and by Assumption 5, we can use the following dual norm inequality (Theorem 6.1  Bühlmann and van de Geer, 2011)
@@
with probability at least $1-\zeta_{t}$. The result of the Theorem follows from choosing $\zeta_{t}=C\mathopen{}\mathclose{{}\left(\sqrt{\text{log}(T)}}\right)^{-1}$, for a large enough constant $C>0$. $\blacksquare$
Appendix C Monte Carlo Experiments

### v3

@@ A.5 Proof of Theorem 3
$\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}v_{t}\mathbf{w}_{t}^{\prime}}}\right|}}\right|_{\infty}=\max\mathopen{}\mathclose{{\left\{\max_{i\leq n_{T}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}v_{t}\widehat{u}_{i,t}}}\right|,\max_{j\leq p_{y}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}v_{t}y_{t-j}}}\right|}}\right\}.$ (9)
@@
Consequently to (9) we have that $Pr(\mathcal{A}_{T})=1-Pr(\mathcal{A}_{T}^{c})\geq 1-Pr\mathopen{}\mathclose{{\left(\max_{i\leq n_{T},l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\widehat{u}_{i,t}}}\right|>\frac{T\lambda}{4}}}\right)-Pr\mathopen{}\mathclose{{\left(\max_{j\leq p_{y},l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}y_{t-j}}}\right|>\frac{T\lambda}{4}}}\right)=1-Pr(I)-Pr(II)$.
We first bound $Pr(I)$. Note that $\widehat{u}_{i,t}=u_{i,t}+x_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}-\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}$. Thus,
$Pr(I)$ $\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\widehat{u}_{i,t}}}\right|>\frac{T\lambda}{4}}}\right)$
@@
$\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|>\frac{T\lambda}{8}}}\right)+\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\mathopen{}\mathclose{{\left(x_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}-\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}}}\right)}}\right|>\frac{T\lambda}{8}}}\right)$ (10)
$=\sum_{i=1}^{n_{T}}Pr(I_{1})+\sum_{i=1}^{n_{T}}Pr(I_{2}).$
We proceed to analyze $I_{1}$. By Assumptions 1, 2 ($b$) and Theorems 17.5, 17.9 and 17.10 in davidson1994, we have that $\{v_{t}u_{i,t}\}$ is an $L_{m}$-mixingale of appropriate size. By the union bound, the Markov’s inequality and the Hansen’s mixingale concentration inequality, it follows that
$Pr\mathopen{}\mathclose{{\left(\max_{i\leq n_{T},l\leq T}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|}}\right]>\frac{T\lambda}{8}}}\right)\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|}}\right]>\frac{T\lambda}{8}}}\right)\leq$
$\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}}\sum_{i=1}^{n_{T}}E\mathopen{}\mathclose{{\left[\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|^{c_{1}}}}\right]\leq\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}}\sum_{i=1}^{n_{T}}C_{1}^{c_{1}}\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}d_{t}^{2}}}\right)^{c_{1}/2}\leq$
$Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}}\ \ .$
For $I_{2}$, note that $v_{t}(x_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}-\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})})=v_{t}\mathopen{}\mathclose{{\left(\sum_{l=1}^{{p}_{i}}(\phi_{i,l}-\widehat{\phi}_{i,l})x_{i,t-l}+\sum_{k=1}^{{q}_{i}}(\theta_{i,k}-\widehat{\theta}_{i,k})u_{i,t-k}}}\right)$. We assume uniform ARMA estimation, namely, there exists $c_{0}>0$ such that $Pr(\mathcal{B}_{T}^{c})\rightarrow 0$, where $\mathcal{B}_{T}\coloneq\mathopen{}\mathclose{{\left\{\max_{1\leq i\leq n_{T}}||\boldsymbol{\widehat{\vartheta}}_{i}-\boldsymbol{\vartheta}_{i}||_{1}\leq c_{0}}}\right\}$, where $\boldsymbol{\vartheta}_{i}=(\phi_{1}\dots,\phi_{p_{i}},\theta_{1},\dots,\theta_{q_{i}})^{\prime}$. Therefore,
$Pr(I_{2})$ $\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{l=1}^{{p}_{i}}(\phi_{i,l}-\widehat{\phi}_{i,l})x_{i,t-l}}}\right|>\frac{T\lambda}{16}}}\right)$
@@
$+\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{k=1}^{{q}_{i}}(\theta_{i,k}-\widehat{\theta}_{i,k})u_{i,t-l}}}\right|>\frac{T\lambda}{16}}}\right)$
$\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{l=1}^{{p}_{i}}x_{i,t-l}}}\right|>\frac{T\lambda}{c_{0}16}}}\right)+Pr\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\sum_{l=1}^{{p}_{i}}\mathopen{}\mathclose{{\left(\phi_{i,l}-\widehat{\phi}_{i,l}}}\right)}}\right|>c_{0}}}\right)+$
$+\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{k=1}^{{q}_{i}}u_{i,t-l}}}\right|>\frac{T\lambda}{c_{0}16}}}\right)+Pr\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\sum_{k=1}^{{q}_{i}}\mathopen{}\mathclose{{\left(\theta_{i,k}-\widehat{\theta}_{i,k}}}\right)}}\right|>c_{0}}}\right)$
$=\sum_{i=1}^{n_{T}}Pr(I_{2}^{\prime})+\sum_{i=1}^{n_{T}}Pr(I_{2}^{\prime\prime})+2Pr(\mathcal{B}_{T}^{c}),\ \ \ \ \ \ \text{for }c_{0}>0.$
Since $\Pr(\mathcal{B}_{T}^{c})=o(1)$, it is sufficient to bound $I_{2}^{\prime}$. Following the same procedure for $I_{1}$, we have
$Pr(I_{2}^{\prime})\leq Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{c_{0}16}}}\right)^{-c_{1}}.$
Therefore, we have that
$Pr(I_{1})\leq Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}},\ \ \ \ \ \ \ Pr(I_{2})\leq Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{c_{0}16}}}\right)^{-c_{1}},$
which implies that
$Pr(I)\leq C_{1}n_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}},$ (11)
for some large enough constant $c_{0}$.
For $Pr(II)$, we follow the same procedure and obtain
$Pr\mathopen{}\mathclose{{\left(\max_{j\leq p_{y},l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}y_{t-j}}}\right|>\frac{T\lambda}{4}}}\right)\leq C2p_{y}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{4}}}\right)^{-c_{1}}.$ (12)
Combining the results from (11) and (12), we obtain that, for $T$ and $n_{T}$ large enough,
$Pr(\mathcal{A}^{c})\leq C(2n_{T}+p_{y})T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{4}}}\right)^{-c_{1}}.$
@@
This means that $\Pr(\mathcal{A}_{T})\geq 1-C\mathopen{}\mathclose{{\left(2n_{T}+p_{y}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{T}\lambda}}}\right)^{c_{1}}$. We impose that the probability of the complement event is bounded by a sequence $\zeta_{T}\rightarrow 0$. Thus, $(2n_{T}+p_{y})(\lambda\sqrt{T})^{-c_{1}}\leq\zeta_{T}$, from which $\lambda\geq\frac{C(2n_{T}+p_{y})^{1/c_{1}}\zeta_{T}^{-1/c_{1}}}{\sqrt{T}}$. The Theorem follows from choosing $\zeta_{t}=C\mathopen{}\mathclose{{\left(\sqrt{\text{log}(T)}}}\right)^{-1}$, for a large enough constant $C>0$. $\square$
A.6 Proof of Theorem 4
The proof of Theorem 4 follows that of Theorem 1 in ADAMEK2023.
Proof of Theorem 4.
By Lemma 6.1 in Buhlmann2011 we obtain
@@ A.6 Proof of Theorem 4
The penalty $\lambda$ is chosen such that $T^{-1}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}\mathbf{v}}}\right|}}\right|_{\infty}\leq\lambda$. Theorem 3, the event $\mathcal{A}_{T}\coloneqq\mathopen{}\mathclose{{\left\{T^{-1}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}\mathbf{v}}}\right|}}\right|_{\infty}\leq\frac{\lambda_{0}}{2}}}\right\}$ holds with high probability, where $\lambda_{0}\leq\frac{\lambda}{2}$. Since $\lambda\geq 2\lambda_{0}$ under $\mathcal{A}_{T}$ and by Assumption 3 of the main text, we can use the following dual norm inequality (Theorem 6.1 Buhlmann2011)
@@
with probability at least $1-\zeta_{t}$. The result of the Theorem follows from choosing $\zeta_{t}=C\mathopen{}\mathclose{{\left(\sqrt{\text{log}(T)}}}\right)^{-1}$, for a large enough constant $C>0$. $\square$
Appendix B Upper Bound for $\psi_{min}$
Here, we would like to point out the role of $\widehat{c}_{ij}^{x}$ for $\widehat{\psi}_{min}^{x}$. To this end, we start by recalling an inequality that links off-diagonal elements and eigenvalues of $\widehat{\mathbf{C}}_{x}$; namely, $\widehat{\psi}_{min}^{x}\leq 1-\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$. Because of this, for any given $\tau\in[0,1)$ we have
which emphasizes how the probability of a generic sample correlation being large in absolute value affects the probability of the minimum eigenvalue being small – and thus the estimation error bounds of the LASSO, as established by Bickel2009. As the next example shows, point the inequality $\widehat{\psi}_{min}^{x}\leq 1-\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$ can be easily fixed.
Example 1:
Let $e_{i}$ and $e_{j}$ be vectors from the standard basis of $\mathbb{R}^{n}$, $i,j\in{1,\dots,N}$. Moreover, let $x_{\pm}=2^{-1/2}(e_{i}\pm e_{j})$, satisfying $x_{\pm}||_{2}=1$, and let $A$ be a correlation matrix with $a_{k}$ be the $k$-th column. Then we have
Thus, $\psi_{min}\leq 1-|a_{ij}|$ for all $i\neq j$ and so
$\psi_{min}\leq 1-\max_{i\neq j}|a_{ij}|\ \ .$
Acknowledgements
Appendix A ARMAr-LS: Simulation Experiments
Consider the univariate regression model
$y_{t}=\alpha x_{t-1}+\varepsilon_{t},\ \ \ \ \ \ t=\,\dots,T,$ (S.13)
where $x_{t}=\phi x_{t-1}+u_{t}$, $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t}$, $u_{t}\sim i.i.d.N(0.\sigma^{2})$, and $\omega_{t}\sim i.i.d(0,\sigma^{2}_{\omega})$. In this section, we provide the estimation and inferential properties of the ARMAr Least Squares (ARMAr-LS) estimator relative to DGP (S.13). In this case, the ARMAr-LS model is
$y_{t}=\alpha u_{t-1}+\phi_{y}y_{t-1}+v_{t},\ \ \ \ \ \ t=\,\dots,T,$ (S.14)
where $v_{t}=(\phi-\phi_{y})x_{t-1}+(\phi_{\varepsilon}-\phi_{y})\varepsilon_{t-1}+\omega_{t}$, and $\phi_{y}=\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}\frac{\phi_{i}\alpha_{i}^{*2}}{1-\phi_{i}^{2}}+\frac{\phi_{\varepsilon}}{1-\phi_{\varepsilon}^{2}}}}\right)/\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}\frac{\alpha_{i}^{*2}}{1-\phi_{i}^{2}}+\frac{1}{1-\phi_{\varepsilon}^{2}}}}\right)$. By Assumption 1, we have $E(u_{t}y_{y-1})=0$, $E(v_{t}|u_{t})=0$, and $E(v_{t}|y_{t-1})=(\phi-\phi_{y})x_{t-1}+(\phi_{\varepsilon}-\phi_{y})\varepsilon_{t-1}\neq 0$ (see Example 2). Consequently, $\widehat{\alpha}=\alpha+\frac{\sum_{t=1}^{T-1}u_{t}v_{t}}{\sum_{t=1}^{T-1}u_{t}^{2}}$ and we have:
  • Unbiasedness. $E(\widehat{\alpha}|\mathbf{u})=\alpha+\frac{\sum_{t=1}^{T-1}u_{t}E(v_{t}|\mathbf{u})}{\sum_{t=1}^{T-1}u_{t}^{2}}=\alpha$;
  • Consistency. $plim(\widehat{\alpha})=\alpha+\frac{E(u_{y}v_{t})}{E(u_{t}^{2})}=\alpha$, consequently to exogeneity of $u_{t}$.
  • Efficiency under common AR(1) restriction. Let $\phi=\phi_{\varepsilon}$, then $v_{t}=\omega_{t}$ and $Var(\widehat{\alpha}|\mathbf{u})=\frac{\sigma_{\omega}^{2}}{\sum_{t=1}^{T-1}u_{t}^{2}}$
We examine, via simulation, the sampling properties of ARMAr-LAS. We explore performances under DGP (S.13) with $T=100,1000$, $\phi=\phi_{\varepsilon}=0.3,0.6,0.9$, and $\sigma^{2}=\sigma_{\omega}^{2}=1$. We compare OLS, OLS augmented with $y_{t-1}$ as an additional regressor (OLSy), and our proposed ARMAr-LS in estimating $\alpha$. For completeness, we also report the estimates obtained from the working model (S.14) (WM). Results are obtained on 1000 Monte Carlo simulations. Figure S.5 illustrates that, as $\phi$ increases, the variance of OLS estimates rises, reflecting their inefficiency under stronger serial correlation, while OLSy becomes increasingly biased. As the sample size $T$ grows, however, all methods exhibit reduced variance. The ARMAr-LASSO estimates outperform those of OLS and OLSy and do not differ significantly from those of the WM, confirming that the estimation of $u$ does not pose any issue.
Figure S.5: Estimates of $\alpha$ on 1000 Monte Carlo simulations for various values of $T$ and $\phi$.
Appendix B Distribution of Sample Correlation Between Serially Correlated Processes: Simulation Experiments